### v1

@@ Abstract
The metabolic response controlled by p53 is mediated through the AMP-activated protein kinase (AMPK), a sensor attuned to cellular metabolic stress conditions [55]. When p53 is activated by AMPK, typically in response to metabolic adversity such as those experienced by cancer cells, p53 restrains cell growth and division, conserves energy, and shifts the cell towards oxidative phosphorylation for more efficient energy production [29]. This can elucidate why cancer cells with p53 mutations tend to rely more on glycolysis and have a higher ability to grow and survive even under stress conditions.
@@
While numerous p53 targets involved in cellular metabolism have been identified, their complex molecular interactions across different scenarios remain largely unexplored. In this study, we leverage the power of mathematical modelling to unveil the intricate machinery behind the regulation of cancer glucose metabolism by p53. By transforming existing experimental data into a mathematical framework, we uncovered hard-to-detect mechanisms and quantitatively analysed the activities of glycolysis and OXPHOS pathways under different cellular states. This methodology provides valuable insights for developing targeted therapeutic approaches aimed at disrupting cancer metabolism and combating the aggressive behaviour of cancer. Although our primary focus is on colon cancer cells, the model’s applicability extends to many cancer types experiencing similar conditions.
2 Model and Assumptions
@@
We developed a comprehensive theoretical framework aiming to delineate the role of p53 on cellular metabolism, particularly its involvement in the Warburg effect. While cancer cells engage in diverse metabolic pathways, the Warburg effect is closely associated with alterations in glucose metabolism. Consequently, our primary attention was devoted to glucose metabolism, investigating its main pathways: glycolysis and oxidative phosphorylation. Integrating information from literature, our model incorporated all well-established p53 targets that markedly manipulate these pathways alongside the signalling pathways commonly activated in cancer in response to growth factors or metabolic stress, influencing the decision-making between these pathways, see Fig. 1. The following section provides an overview of the cellular events and the molecule interactions considered while constructing our model.
2.1 Model assumptions
2.1.1 Growth factors activate PI3K/AKT pathway
2.1.2 AKT mediates mTOR activating
Upon activation of AKT, various downstream substrates are phosphorylated. One critical effector of AKT is the mammalian target of rapamycin (mTOR) [22, 49]. The activation of mTORC1 is vital for the control of cellular processes such as cell growth and metabolism, primarily through its ability to regulate the mRNA translation [22, 49, 26].
2.1.3 mTOR induces the expression of HIF1
HIF1 is a transcription factor consisting of two subunits, HIF1 and HIF1. The -subunit is constitutively present in the nucleus, while the -subunit is rarely detectable in normoxia but strikingly induced in hypoxic conditions [115, 75, 35, 78]. The HIF1 gene is continuously transcribed and translated, but its expression stays low due to rapid destruction via the ubiquitin-proteasome pathway in an oxygen-dependent manner [115, 75, 35].
2.1.4 HIF1 promotes glycolysis pathway
@@
2.1.5 Metabolic stress activates AMPK
2.1.6 AMPK activates tumour suppressor gene p53
2.1.7 p53 induces genes negatively regulate its activation
[…]
2.1.8 p53 induces genes to promote its activation, simultaneously inhibiting HIF1
[…]
2.1.9 p53 suppresses glycolysis and enhances mitochondrial respiration
[…]
2.2 Model reactions
2.3 Model equations
Our model network spans three compartments—cytoplasm, nucleus, and mitochondria—where directional fluxes determine the temporal change in concentration of 33 molecules within the glucose oxidation pathways, schematically shown in Fig. 1. Consequently, our system is entirely governed by 33 differential equations where cytoplasmic, nuclear, and mitochondrial molecular species are represented in our equations by the subscripts ‘c’, ‘n’, and ‘m’, respectively, while the ‘’ superscript symbol indicates active species for those existing in both active and inactive states.
Each term in the model represented by ’v’ notation corresponds to one of the reactions detailed in the model reaction section, where or . Despite the model’s complexity, involving 67 parameters, it demonstrates significant robustness, showing minimal sensitivity to only a few of these parameters (Table 11). For more details, the comprehensive model incorporating explicit differential equations, corresponding chemical reactions, parameter values, parameter experimental justifications, and model robustness analysis is provided in the Appendix section (A). Numerical simulations for this system were conducted using the ’ode’ routine in MATLAB and Gear’s method in the XPPAUT software.
@@ 3 Results
Motivated by these findings, we studied the impact of inhibiting mTOR activity on metabolic pathways and energy production levels under hypoxic conditions. We constructed a bifurcation diagram showcasing the metabolic activity of glycolysis and OXPHOS and their contributions to energy production under different k35 rates (mTOR-dependent HIF1 synthesis rate), see Fig. 4. Analysing these diagrams confirms the mTOR involvement in producing high energy levels in hypoxic cancer cells, as impeding its activity drove the cell towards a similar energy level produced in our hypoxic normal cells.
@@ 3.4 Dual stable steady states in cancer cells, contrasted by singular stability in normal cells
The transition between these two states in p53 wild-type cells is governed by the phosphorylation levels of AMPK, the protein responsible for instigating the p53-metabolic stress response. This dynamic is further elucidated by the bifurcation diagram, illustrating the levels of p53 and HIF1 under various AMPK phosphorylation rates denoted as k24 (Fig. 6).
Under low phosphorylation rates of AMPK, cells exhibit two stable steady states: high activation of p53 (Stable SS p53+/+) and no activation of p53 (Stable SS p53-/-). However, exceeding the bifurcation point by increasing the phosphorylation rate (k24) induces p53 activation and shifts the cell into a unique stability regime, representing the p53-wild-type state.
@@ 3.5 Restoring normal metabolism in cancer cells by increasing the p53 activation levels
Inspired by these insights, we delved into the impact of boosting SCO2 levels on the metabolic phenotypes of cancer cells, particularly those with p53 mutations. In-silico, we elevated the SCO2 expression levels of p53-/- cells by increasing its basal production rate, k18 (Fig. 9).
@@ 4 Discussion
In the context of existing research, previous computational modelling sheds light on different aspects of cancer metabolism —ranging from the effect of reactive oxygen species (ROS) on HIF1 stabilisation in ischemic conditions [91] to the identification of metabolic targets to hinder cancer migration [126]. Despite these insights, the genetic complexities underpinning the Warburg effect remain elusive. A notable contribution in this domain has been introduced recently by Linglin et al., discussing the genetic regulation of the interplay between glycolysis and oxidative phosphorylation [127]. However, this work did not take the vital influence of p53 into account and did not perform a quantitative analysis of how genetic factors impact metabolic outcomes or potential strategies to mitigate the Warburg effect. Our research bridges this gap by constructing a gene-based mathematical framework that dissects the mechanisms through which p53, alongside other genetic regulations, influences glycolysis and OXPHOS and quantitatively explores their impact on these pathways under various cellular conditions.
@@ Appendix A Detailed Model
A.1 Model descriptions
A.1.1 Cytoplasmic and nuclear p53 equations
In response to metabolic stress encountered by cancer cells, the AMP-activated protein kinase (AMPK) is activated through phosphorylation [37, 43, 31, 74], which then phosphorylates cytoplasmic p53, triggering its stabilisation and nuclear translocation [55, 50]. The breakdown of the p53 protein involves two pathways: a basal degradation and a process mediated by murine double minute 2 (MDM2) via the ubiquitin-proteasome pathway [89, 42]. The dephosphorylation of nuclear p53 is catalysed by wild-type p53-induced phosphatase 1 (WIP1) [12], which unmasks p53 nuclear export signals, thus directing them back to the cytoplasm. Let the concentration of cytoplasmic p53, nuclear p53, active cytoplasmic AMPK, cytoplasmic MDM2, nuclear MDM2, and nuclear WIP1 be denoted by , , , , , and , respectively.
@@ A.1.1 Cytoplasmic and nuclear p53 equations
A.1.2 Cytoplasmic and nuclear MDM2 equations
Upon activation, tetrameric nuclear p53 triggers the production of its negative regulator, MDM2, in the cytoplasm [9], a process modelled by a Hill function with a coefficient four. Although MDM2 predominantly resides within the cytoplasm [81], stimulation by growth factor signals causes protein kinase B (AKT) to phosphorylate MDM2, facilitating its relocation to the nucleus to hinder p53 transcriptional activity [79, 124]. Let denote the active cytoplasmic AKT concentration.
@@ A.1.2 Cytoplasmic and nuclear MDM2 equations
A.1.3 WIP1, PTEN, SCO2, and TIGAR equations
WIP1, phosphatase and tensin homolog (PTEN), synthesis of cytochrome c oxidase 2 (SCO2), and TP53-inducible glycolysis and apoptosis regulator (TIGAR) are all targets of p53, synthesised under normal conditions. However, with the activation of p53, nuclear p53 instigates the activation of their promoters, enhancing their production rate [12, 29, 84, 13, 67, 103]. Among these, WIP1 forms a negative feedback loop that inhibits p53 function, whereas PTEN contributes to a positive feedback loop that supports p53 activity. Let the concentration of the cytoplasmic PTEN, mitochondrial SCO2, and cytoplasmic TIGAR be denoted by , , and , respectively.
@@ A.1.3 WIP1, PTEN, SCO2, and TIGAR equations
A.1.4 Active AMPK, PIP3, AKT, and mTOR equations
Along with AMPK activation by metabolic stress, p53 also triggers AMPK activation through its targets, sestrin1 and 2, which interact with the -subunits of AMPK, resulting in its phosphorylation [10, 101]. Our model incorporates this process by assuming a direct activation of AMPK by nuclear p53, forming a robust positive feedback loop. On the other hand, growth factor stimulation leads to the phosphorylation of phosphatidylinositol 4,5-bisphosphate (PIP2) into phosphatidylinositol 3,4,5-trisphosphate (PIP3), facilitating AKT phosphorylation by biding to PIP3 [21, 114, 18]. This series of events activates AKT, which then influences the mammalian target of rapamycin (mTOR) activation [22, 49]. However, upon p53 activation, p53 target PTEN converts PIP3 back to PIP2, effectively attenuating the growth factors signalling pathway [80, 18, 29]. Furthermore, the activation of AMPK acts as a regulatory mechanism that suppresses mTOR activity to constrain the protein synthesis process induced by mTOR [10, 101, 29]. Let the concentration of cytoplasmic PIP3 and active cytoplasmic mTOR be denoted by and , respectively, while , , and represent the concentrations for the inactive form of the corresponding variable.

### v2

@@ Abstract
The metabolic response controlled by p53 is mediated through the AMP-activated protein kinase (AMPK), a sensor attuned to cellular metabolic stress conditions [55]. When activated by AMPK, typically in response to metabolic adversity such as those experienced by cancer cells, p53 restrains cell growth and division, conserves energy, and shifts the cell towards oxidative phosphorylation for more efficient energy production [29]. This may explain why cancer cells with p53 mutations tend to rely more on glycolysis and have a higher ability to grow and survive even under stress conditions.
Previous mathematical modelling sheds light on different aspects of cancer metabolism —ranging from the effect of reactive oxygen species (ROS) on HIF1 stabilisation in ischemic conditions [93] to the identification of metabolic targets to hinder cancer migration [128]. Despite these insights, the genetic complexities underpinning the Warburg effect remain elusive. Recently, Linglin et al. made a notable contribution by discussing the genetic regulation of the interplay between glycolysis and oxidative phosphorylation [129]. Nonetheless, this work did not account for the crucial influence of p53 or demonstrate its impact on metabolic pathways as observed experimentally in cancer cells.
Undoubtedly, the tumour-suppressive role of p53 has been studied through several computational modelling approaches. Ma et al. modelled the oscillatory dynamics of p53, emphasising the correlation between the DNA damage severity and the average number of p53 pulses [90]. Further studies by Zhang et al. have elucidated how the frequency of p53 pulses can influence cellular decisions between repair and apoptosis [132] and described the transition from cell repair to apoptosis with an inverse relationship between the apoptosis time and damage strength [133]. Another study has introduced a two-phase dynamic of p53, distinguishing between partial and complete activation of p53 [137]. Our recent study expanded on these findings by demonstrating that p53 can switch among three dynamic modes in a DNA damage strength-dependent manner following chemotherapy [6]. Despite these advances, earlier works have mainly concentrated on p53 dynamic behaviour in response to DNA damage stimuli and the subsequent cell fate, leaving a substantial gap in understanding its mechanisms regarding cellular metabolism.
@@
To address this gap, this study develops a mathematical model that unveils the intricate machinery behind the regulation of cancer glucose metabolism by p53. While numerous p53 targets involved in cellular metabolism have been identified, their complex molecular interactions across different scenarios remain largely unexplored. By transforming existing experimental data into a mathematical framework, we uncovered hard-to-detect mechanisms and quantitatively analysed the activities of glycolysis and OXPHOS pathways under different cellular states. This methodology provides valuable insights for developing targeted therapeutic approaches aimed at disrupting cancer metabolism and combating the aggressive behaviour of cancer. Although our primary focus is on colon cancer cells, the model’s applicability extends to many cancer types experiencing similar conditions.
2 Methods
@@
We constructed a comprehensive theoretical framework aiming to delineate the role of p53 on cellular metabolism, particularly its involvement in the Warburg effect. While cancer cells engage in diverse metabolic pathways, the Warburg effect is closely associated with alterations in glucose metabolism. Consequently, our primary attention was devoted to glucose metabolism, investigating its main pathways: glycolysis and oxidative phosphorylation. Integrating information from literature, our model incorporated all well-established p53 targets that markedly manipulate these pathways alongside the signalling pathways commonly activated in cancer in response to growth factors or metabolic stress, influencing the decision-making between these pathways, see Fig. 1. The following section briefly discusses these signalling pathways and their involvement in glucose metabolism.
2.1 Signalling pathways in glucose metabolism
2.1.1 Growth factor signalling pathway
In response to growth factors, extracellular molecules activate specific receptors on the cell membrane, initiating the intracellular activation of phosphoinositide 3-kinase (PI3K). PI3K then phosphorylates phosphatidylinositol 4,5-bisphosphate (PIP2) into phosphatidylinositol 3,4,5-trisphosphate (PIP3), facilitating the protein kinase B (AKT) activation [21, 116, 18].
Upon AKT activation, a cascade of events occurs, ultimately activating the mechanistic target of rapamycin (mTOR) [22, 49], which subsequently promotes the synthesis of proteins required for cell growth and proliferation. One of these proteins is HIF1 [26, 75, 41, 111], a transcription factor that drives the expression of genes essential for glycolytic metabolism. These genes include glucose transporter-1 and -3 (GLUT1) and (GLUT3), respectively [1], lactate dehydrogenase (LDH) [117], and pyruvate dehydrogenase kinase-1 and -3 (PDK1) and (PDK3), respectively [4, 68, 122, 62].
2.1.2 Metabolic stress signalling pathway
Different forms of metabolic stress activate AMPK, a crucial enzyme that restores cellular energy balance by suppressing ATP-consuming processes and promoting ATP production [37, 43, 31, 74]. Through this activation, AMPK restrains cellular growth and increases ATP production efficiency primarily by repressing mTOR activation and stimulating the transcriptional activity of p53 [31, 74, 51].
2.1.3 p53 transcriptional targets and feedback mechanisms
p53 exerts its effects through the transcriptional regulation of a wide array of target genes involved in various cellular processes. Some p53 targets form a negative feedback loop that dampens its stability, such as MDM2 and wild-type p53-induced phosphatase 1 (WIP1) [9, 12]. WIP1 dephosphorylates p53, increasing its susceptibility to MDM2-mediated degradation [12].
Conversely, p53 activates genes that promote its activation and simultaneously inhibit the glycolytic regulator HIF1. For instance, p53 induces the expression of proteins that activate AMPK, which in turn inhibits mTOR activity and its downstream target HIF1 [10, 103, 29]. This process forms a positive feedback loop, as active AMPK phosphorylates p53, further enhancing its stabilization [29]. Another crucial target of p53 is phosphatase and tensin homolog (PTEN) [105], which attenuates the PI3K/AKT/mTOR pathway by converting PIP3 back to PIP2 [80, 18, 29]. The induction of PTEN also supports p53 by hindering the AKT-dependent translocation of MDM2 to the nucleus, thereby boosting the transcriptional activity of p53 [80, 89].
In addition, p53 regulates genes directly involved in glucose metabolic pathways. This includes diminishing the protein synthesis of GLUT1, GLUT3 [1, 101, 56], and PDK2 [4, 66], while stimulating the expression of other molecules, such as the TP53-inducible glycolysis and apoptosis regulator (TIGAR) [13, 67], and the synthesis of cytochrome c oxidase 2 (SCO2) [84].
2.1.4 Glucose metabolism pathways
Glucose metabolism comprises three main stages: starting with glycolysis, advancing to the tricarboxylic acid (TCA) cycle, and finishing with the electron transport chain (ETC).
@@ 2.1.4 Glucose metabolism pathways
Glucose is transported to the cells by specific glucose transporters located in the cell membrane [101, 108, 1, 83], namely GLUT1 and GLUT3 in our model, which are negatively regulated by p53 and positively by HIF1, controlling glucose uptake rate [1, 101, 56]. Once inside the cell, glucose undergoes an irreversible conversion to glucose-6-phosphate (G6P), consuming one ATP molecule [35]. G6P can then either proceed through glycolysis or enter the pentose phosphate pathway (PPP) [53], a decision influenced by the p53 target TIGAR, which inhibits the enzyme catalysing the third step in the glycolysis pathway, diminishing glycolysis flux [13, 67]. Continuing with glycolysis, a molecule of G6P is converted into two molecules of pyruvate, yielding three net ATP and two reduced nicotinamide adenine dinucleotide (NADH) molecules [117, 35].
In the presence of oxygen, pyruvate derived from glycolysis typically enters the mitochondria for further energy production through oxidative phosphorylation. However, increased levels of LDH enzyme, driven by HIF1 activation, redirect pyruvate away from the mitochondria by catalysing its conversion to lactate and oxidising NADH back to NAD+ [117, 35]. This shift allows cells to produce ATP less efficiently through glycolysis while consuming more glucose [117].
@@
Within the mitochondria, the pyruvate dehydrogenase (PDH) complex facilitates the first step towards glucose respiration by catalysing the oxidative decarboxylation of pyruvate to acetyl-CoA, concurrently reducing one NAD+ into NADH [122, 121, 94]. Acetyl-CoA subsequently enters the TCA cycle, undergoing a series of chemical reactions that generate one energy molecule, three NADH, and one reduced flavin adenine dinucleotide (FADH2) molecule [88].
The function of the PDH enzyme is controlled by the PDK family, which phosphorylates the E1- subunit of PDH, blocking its decarboxylation activity [122, 121, 94]. HIF1 enhances the promoter activities of two PDK family members, namely PDK1 and PDK3 [4, 68, 122, 62], while p53 suppresses the expression of another PDK member called PDK2 [4, 66].
@@
In the last stage, NADH and FADH2 are oxidised back into NAD+ and FAD via protein complexes in the inner mitochondrial membrane [7]. This oxidation process involves electron transfer from NADH and FADH2 across these complexes, during which protons are pumped from the mitochondrial matrix to the intermembrane space. This action creates a proton gradient that drives protons back into the matrix, facilitating ATP production [7].
Among these complexes, Complex IV acts as the final electron acceptor, channelling the electrons to molecular oxygen by using 0.5 oxygen (O2) molecules for each pair of electrons received from NADH or FADH2 [7]. However, the activity of Complex IV is regulated by SCO2, a p53-regulated gene essential for its proper assembly and maturation. Thus, deficiency of SCO2 can impair Complex IV functionality and disrupt the electron flow within the ETC [84, 120].
2.2 Mathematical framework
Our model encompasses the entire glucose oxidation network, spanning three compartments—cytoplasm, nucleus, and mitochondria. The directional fluxes within these compartments drive the temporal changes in the concentrations of 33 molecules involved in glucose oxidation pathways, as illustrated in Fig. 1. Consequently, the system is governed by 33 differential equations where cytoplasmic, nuclear, and mitochondrial molecular species are represented by the subscripts ‘c’, ‘n’, and ‘m’, respectively, while the ‘’ superscript symbol indicates active species for those existing in both active and inactive states. Each term in the model represented by ’v’ notation corresponds to one of the reactions detailed in the model reactions section in the Appendix (A), where or .
This model applies to all cell types, both normal and cancerous, with certain restrictions. Under the assumption that normal cells exist in a healthy environment, metabolic stress and continuous activation of growth factor signals are exclusively attributed to cancer cells. Consequently, and are set to zero in normal cells. Additionally, to differentiate between cancer cell types, reducing to zero in p53-mutated cancer cells can ensure the absence of p53 response in these cells.
For a comprehensive understanding, detailed descriptions of the model—including assumptions, reactions, parameter values, experimental justifications, and sensitivity analysis—are provided in the Appendix (A). Additionally, all numerical simulations in this study were performed using MATLAB’s ’ode’ routine and Gear’s method in the XPPAUT software.
@@ 3 Results
Motivated by these findings, we studied the impact of inhibiting mTOR activity on metabolic pathways and energy production levels under hypoxic conditions. We constructed diagrams showcasing the metabolic activity of glycolysis and OXPHOS and their contributions to energy production under different rates (mTOR-dependent HIF1 synthesis rate), see Fig. 4. Analysing these diagrams confirms the mTOR involvement in producing high energy levels in hypoxic cancer cells, as impeding its activity drove the cell towards a similar energy level produced in our hypoxic normal cells.
@@ 3.4 Dual stable steady states in cancer cells, contrasted by singular stability in normal cells
The transition between these two states in p53 wild-type cells is governed by the phosphorylation levels of AMPK, the protein responsible for instigating the p53-metabolic stress response. This dynamic is further elucidated by the bifurcation diagram, illustrating the levels of p53 and HIF1 under various AMPK phosphorylation rates denoted as (Fig. 6).
Under low phosphorylation rates of AMPK, cells exhibit two stable steady states: high activation of p53 (Stable SS p53+/+) and no activation of p53 (Stable SS p53-/-). However, exceeding the bifurcation point by increasing the phosphorylation rate () induces p53 activation and shifts the cell into a unique stability regime, representing the p53-wild-type state.
@@ 3.5 Restoring normal metabolism in cancer cells by increasing the p53 activation levels
Inspired by these insights, we delved into the impact of boosting SCO2 levels on the metabolic phenotypes of cancer cells, particularly those with p53 mutations. In-silico, we elevated the SCO2 expression levels of p53-/- cells by increasing its basal production rate, (Fig. 9).
@@ 4 Discussion
In terms of existing research, Linglin et al. have made good strides in elucidating the influence of genetic regulation on glycolysis and oxidative phosphorylation, identifying a hybrid metabolic state unique to cancer cells [129]. However, this study did not consider the vital influence of p53, nor did it conduct a quantitative analysis of how genetic factors impact metabolic outcomes or potential strategies to mitigate the Warburg effect. Our research bridges these gaps by constructing a comprehensive mathematical framework that dissects the mechanisms through which p53, alongside other genetic regulations, influences glycolysis and OXPHOS and quantitatively explores their impact on these pathways under various cellular conditions.
@@ Appendix A Detailed Model
A.1 Model assumptions
This section provides a detailed description of the cellular events and the molecule interactions considered while constructing our model.
A.1.1 Growth factors activate PI3K/AKT pathway
A.1.2 AKT mediates mTOR activating
Upon activation of AKT, various downstream substrates are phosphorylated. One critical effector of AKT is the mechanistic target of rapamycin (mTOR) [22, 49]. The activation of mTORC1 is vital for the control of cellular processes such as cell growth and metabolism, primarily through its ability to regulate the mRNA translation [22, 49, 26].
A.1.3 mTOR induces the expression of HIF1
Hypoxia-inducible factor 1 (HIF1) is a transcription factor consisting of two subunits, HIF1 and HIF1. The -subunit is constitutively present in the nucleus, while the -subunit is rarely detectable in normoxia but strikingly induced in hypoxic conditions [117, 75, 35, 78]. The HIF1 gene is continuously transcribed and translated, but its expression stays low due to rapid destruction via the ubiquitin-proteasome pathway in an oxygen-dependent manner [117, 75, 35].
A.1.4 HIF1 promotes glycolysis pathway
[…]
A.1.5 Metabolic stress activates AMPK
A.1.6 AMPK activates tumour suppressor gene p53
A.1.7 p53 induces genes negatively regulate its activation
[…]
A.1.8 p53 induces genes to promote its activation, simultaneously inhibiting HIF1
[…]
A.1.9 p53 suppresses glycolysis and enhances mitochondrial respiration
[…]
A.2 Model reactions
A.3 Model equations
The constructed differential equations, alongside their corresponding chemical reactions and parameter values, are explicitly presented below.
A.3.1 Cytoplasmic and nuclear p53 equations
Let the concentration of cytoplasmic p53, nuclear p53, active cytoplasmic AMPK, cytoplasmic MDM2, nuclear MDM2, and nuclear WIP1 be denoted by , , , , , and , respectively.
@@ A.3.1 Cytoplasmic and nuclear p53 equations
A.3.2 Cytoplasmic and nuclear MDM2 equations
Let denote the active cytoplasmic AKT concentration.
@@ A.3.2 Cytoplasmic and nuclear MDM2 equations
A.3.3 WIP1, PTEN, SCO2, and TIGAR equations
Let the concentration of the cytoplasmic PTEN, mitochondrial SCO2, and cytoplasmic TIGAR be denoted by , , and , respectively.
@@ A.3.3 WIP1, PTEN, SCO2, and TIGAR equations
A.3.4 Active AMPK, PIP3, AKT, and mTOR equations
Let the concentration of cytoplasmic PIP3 and active cytoplasmic mTOR be denoted by and , respectively, while , , and represent the concentrations for the inactive form of the corresponding variable.